OpenStudy (anonymous):

completing the square-- 3x^2+6x-5=0

OpenStudy (anonymous):

\[=3(x ^{2}+2x)-5\] \[=3(x ^{2}+2x +1-1)-5\] \[=3(x+1)^{2}-3-5\] \[=3(x+1)^{2}-8\]